What the x and the y in substitution system of equations. 3x+y=9 y=-5x+9 show details please.
Answer by jim_thompson5910(35256)

3x+y=9

3x+(-5x+9)=9

3x-5x+9=9

-2x+9=9

-2x=9-9

-2x=0

x=0/(-2)

x=0

--------------------------
y = -5x+9

y = -5(0)+9

y = 0+9

y = 9


So the solutions are x = 0 and y = 9
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------